When multiply 6/13 by inverse of 7/13 we get​
pliiiiiiiiiiiiizzzzz answer it clearly

Grade:12

1 Answers

SJ
3 years ago
welcome to Askiitians

inverse of 7/13= 13/7

so their multiplication =6/13*13/7
= 6/7
ans

Thanks